OpenStudy (hannaha):

MEDAL FOR HELP!!! Four numbers have a total of 92. The first number is twice the second number and the third number is 4 more than the sum of the first two. The fourth number is 3 less than the difference of the first two numbers. What are the numbers? The first number is ___ , the second number is ___ , the third number is ____ , and the fourth number is ____ .

OpenStudy (valerievoun):

a = 26 b = 13 c = 43 d = 10

OpenStudy (hannaha):

ok but how do you get that?

OpenStudy (valerievoun):

a+b+c+d = 92 "The four numbers have a total of 92" 2) a = 2b "The first number is twice the second number..." 3) c = (a+b)+4 "...the third number is 4 more than the sum of the first two." 4) d = (a-b)-3 "The fourth number is 3 less than the difference of the first two numbers." Substitute equation 2) into equation 1) to get: 1a) 2b+b+c+d = 92 Simplify. 2a) 3b+c+d = 92 Now substitute equation 3 into this equation. 2b) 3b+(a+b)+4+d = 92 Simplify this. 2b) a+4b+d+4 = 92 Substitute equation 2) into this equation. 2c) 2b+4b+d+4 = 94 Simplify. 2c) 6b+d+4 = 92 Substitute equation 4) into this equation. 2d) 6b+((a-b)-3)+4 = 92 Substitute equation 2) into this equation. 2e) 6b+((2b-b)-3)+4 = 92 Simplify. 2e) 7b+1 = 92 Subtract 1 from both sides. 7b = 91 Divide both sides by 7. b = 13 a = 2b = 2(13) = 26 c = (a+b)+4 = 26+13+4 = 43 d = (a-b)-3 = 26-13-3 = 10

OpenStudy (valerievoun):

do you understand?

OpenStudy (hannaha):

hold on I'm trying....hmm

OpenStudy (valerievoun):

don't plagiarize tho if this is an essay lol

OpenStudy (hannaha):

okay this is super confusing -_- sorry

OpenStudy (hannaha):

lol no way this isn't an essay

OpenStudy (valerievoun):

read one step at a time

OpenStudy (hannaha):

so this is what I'm getting a= x b= 2x c= 2x + x + 4 d= 2x - x - 4 is this right? if so now how do i continue?

OpenStudy (hannaha):

@valerievoun

OpenStudy (dayakar):

let the 4 numbers are a,b,c,d a+b+c+d = 92 --------(1) a= 2b-------(2) c= a+b+4 -----(3) d= a-b-3 ----(4)

OpenStudy (hannaha):

right, right ok now what

OpenStudy (dayakar):

plug (2) in (3) and (4) c= 3b+4------(5) d= b-3------(6)

OpenStudy (dayakar):

is it clear

OpenStudy (dayakar):

do u understand

OpenStudy (hannaha):

so what numbers are you adding?

OpenStudy (hannaha):

i'm trying but this is so confusing :(

OpenStudy (dayakar):

in (3) c= a+b+4 c= 2b +b +4 [ from (2) a =2b ] c= 3b +4 ----we are taking it as (5) same way (6)

OpenStudy (dayakar):

i explained above how we get (5)

OpenStudy (dayakar):

if u understood that we go next step

OpenStudy (valerievoun):

@dayakar will guide you through it

OpenStudy (dayakar):

respond quickly , do u get it or not

OpenStudy (dayakar):

now plug (2) ,(5), and (6) in (1) a+b+c+d= 92-----(1) 2b +b +3b+4 +b-3 = 92 7b+1 =92 7b= 91 b= 91/7 b= 13

OpenStudy (dr_feelz):

You can put this solution on YOUR website! Let the four numbers be a, b, c, and d. From the problem description you can write: 1) a+b+c+d = 92 "The four numbers have a total of 92" 2) a = 2b "The first number is twice the second number..." 3) c = (a+b)+4 "...the third number is 4 more than the sum of the first two." 4) d = (a-b)-3 "The fourth number is 3 less than the difference of the first two numbers." Substitute equation 2) into equation 1) to get: 1a) 2b+b+c+d = 92 Simplify. 2a) 3b+c+d = 92 Now substitute equation 3 into this equation. 2b) 3b+(a+b)+4+d = 92 Simplify this. 2b) a+4b+d+4 = 92 Substitute equation 2) into this equation. 2c) 2b+4b+d+4 = 94 Simplify. 2c) 6b+d+4 = 92 Substitute equation 4) into this equation. 2d) 6b+((a-b)-3)+4 = 92 Substitute equation 2) into this equation. 2e) 6b+((2b-b)-3)+4 = 92 Simplify. 2e) 7b+1 = 92 Subtract 1 from both sides. 7b = 91 Divide both sides by 7. b = 13 a = 2b = 2(13) = 26 c = (a+b)+4 = 26+13+4 = 43 d = (a-b)-3 = 26-13-3 = 10 The numbers are: a = 26 b = 13 c = 43 d = 10